# Humans as a Calibration Pattern: Dynamic 3D Scene Reconstruction from Unsynchronized and Uncalibrated Videos

### v1

@@ Abstract
Recent works on dynamic neural field reconstruction assume input from synchronized multi-view videos with known poses. These input constraints are often unmet in real-world setups, making the approach impractical. We demonstrate that unsynchronized videos with unknown poses can generate dynamic neural fields if the videos capture human motion. Humans are one of the most common dynamic subjects whose poses can be estimated using state-of-the-art methods. While noisy, the estimated human shape and pose parameters provide a decent initialization for the highly non-convex and under-constrained problem of training a consistent dynamic neural representation. Given the sequences of pose and shape of humans, we estimate the time offsets between videos, followed by camera pose estimations by analyzing 3D joint locations. Then, we train dynamic NeRF employing multiresolution grids while simultaneously refining both time offsets and camera poses. The setup still involves optimizing many parameters, therefore, we introduce a robust progressive learning strategy to stabilize the process. Experiments show that our approach achieves accurate spatiotemporal calibration and high-quality scene reconstruction in challenging conditions. Project page
@@ 1 Introduction
Recent advances in neural radiance fields (NeRF) have been extended into obtaining dynamic scenes of any topology [30] with small memory consumption [5, 8] using video inputs. However, even with a series of regularization techniques, obtaining a dynamic 3D volume with video inputs is a highly under-constrained problem and often suffers from instability and convergence problems. The claimed success of dynamic NeRF assumes either unrealistic movements of a monocular camera that can emulate a multi-view setup [9] or perfectly calibrated and synchronized multi-view inputs. Such input observations cannot be achieved in casual setups. Time synchronization often relies on hard-wired devices [14, 10] or needs additional cues [6] (e.g., sound peak). Pose estimation methods struggle in scenes with textureless area or repetitive structure, as shown in Fig. 2. Furthermore, it becomes more challenging if scenes are dynamic and videos are not synchronized.
@@
We aim to allow training dynamic NeRF from a set of casually captured real-world videos of a shared event, such as a sports game or a concert, which only can be collected later. In other words, we reconstruct photorealistic 4D scenes from unsynchronized multi-view videos with unknown camera poses. We need strong priors to tackle this challenging problem with high degrees of freedom. We focus on humans, one of the most common dynamic objects in scenes. Human pose estimation in computer vision has progressed rapidly and now demonstrates a reliable performance even in general images or videos. We consider an estimated human parameter a robust mid-level representation to deduce the relationship between unsynchronized videos. Namely, we exploit human motion as a calibration pattern to estimate both time offsets and camera poses.
@@
We first find the time offsets using the sequence of human shape and pose parameters estimated from individual videos. We can obtain a robust estimation of time offsets by enforcing global consistency in the sequence of motions with exhaustive pairwise scores. Specifically, we consider the detected parametric model of a human as a time series whose feature vector is the pose and shape parameters. We apply dynamic time warping (DTW) between every pair of videos to compute a constant offset to minimize the overall discrepancy in the explicit 3D joint positions and record the associated cost value. The pairwise values are stored within a matrix, from which we find a global sequence alignment with the overall minimum costs.
Even with slight misalignments, NeRF results can degrade significantly. Previous works propose optimizing camera poses jointly with the static scene, which only works with a good initialization point. Our initialization comes from the alignment of human motions, by applying Procrustes analysis on aggregated 3D joint positions across all frames. Note that we estimate transformation that maps camera coordinate to coherent global coordinate, so that we can estimate camera poses for individual time steps, making our setup generalizable to moving hand-held cameras.
@@
We then train dynamic NeRF [8] while refining the estimated time offsets and camera poses. We modify coarse-to-fine registration [21] to stabilize the training of the multiresolution grid-based representation. Furthermore, we propose a curriculum learning strategy that adds more variables as the optimization progresses. Optimization scheduling is critical for convergence in gradient-based optimization.
@@
We evaluate our approach in the real-world CMU Panoptic Studio dataset [14] consisting of 9 seconds-long unsynchronized multi-view videos with a maximum of 4 seconds offsets without given camera poses. Experiments show that our initialization stage robustly aligns human motion, followed by a stable refinement stage that results in highly accurate calibration (rotation error $<0.4^{\circ}$, translation error $<0.2$ cm, time offset error $<0.03$ frame in average).
@@ 2 Related Works
Since humans are one of the most common and important objects in the scene, there exist abundant datasets containing humans [13, 25, 1, 22, 10]. Recent methods leverage powerful data-driven prior to estimate human pose and shape from images or videos. Most approaches recover human pose and shape by estimating parametric models such as SMPL [24]. From optimization-based methods [3, 32], which leverage robust 2D joint estimation networks, to regression-based methods [15, 16, 18] that utilize 3D human datasets, early works faithfully recover 3D human pose and shape in camera coordinate. Recent advances [19, 37] take a step forward to estimate the global trajectories of humans, which is crucial for a complete understanding of human motion. Notably, SLAHMR [44] utilizes human motion priors to recover both real-world scaled global human motion and camera trajectory, whose scale is ambiguous with the SLAM pipeline alone. We note that monocular human shape and motion estimation works have matured and generalize well to novel scenes. In this paper, we explore the potential of human motion as a robust mid-level representation for calibration in both temporal and spatial domain.
@@
The emergence of NeRF [28] and its extension to the spatiotemporal domain enables immersive experience in a dynamic world, which was previously limited to a small navigatable area and requires a complicated capturing system [4]. One of the most common approaches to reconstruct 4D scenes is reconstructing 3D scenes in a canonical frame and optimize a deformation field to warp them [29, 30, 33, 20, 23, 43]. However, as pointed out by Park et al. [31], it is hard to let the deformation network learn general dynamics in practice. On the other hand, recent grid-based methods [8, 5] simply represent 4D scenes by introducing additional temporal domain without deformation field, which enables to represent general dynamic scenes easily. We exploit K-Planes [8] for our 4D scene representation due to its versatility.
However, all of the aforementioned dynamic NeRFs need synchronized videos and ground-truth camera poses. Recently, Sync-NeRF [17] aims to reconstruct dynamic NeRF from unsynchronized videos. They additionally optimize time offset parameters during the dynamic NeRF reconstruction process. However, Sync-NeRF requires initial time offsets close to ground truth since it can only deal with small temporal perturbations and they also require known camera poses.
@@
Since iNeRF [45] has shown that one can optimize camera poses given pretrained NeRF with photometric loss, recent works attempt to reconstruct NeRF from unknown camera poses. They ease the burden of the hard requirements of NeRF (i.e., perfect camera poses) by jointly optimizing camera poses during the optimization process of NeRF. Instead of naïvely optimizing parameters [42], recent works show that coarse-to-fine optimization by progressively adding higher frequency bands [21] or curriculum learning strategy that progressively adds camera parameters [11]. However, previous works only optimize camera poses with static NeRF. To the best of our knowledge, we are the first to reconstruct dynamic NeRF from unknown time offsets and camera poses. We also observe that progressive training is critical for robust optimization in dynamic NeRF scenarios as well. We propose a similar coarse-to-fine optimization technique suited for multiresolution grid representation and curriculum learning strategy. More importantly, all previous works need good initial points and only adjust small misalignments during the optimization process or only work for image sets with extremely small baselines such as forward-facing captures [27]. We propose a novel way to obtain good initial points with humans, both for time offsets and camera poses.
@@ 3 Method
We reconstruct dynamic NeRF from unsynchronized multi-view videos with unknown camera poses. We first extract human motion of individual video in its own camera coordinate (Sec. 3.1). We then utilize extracted human motion as a calibration pattern to estimate both time offsets and camera poses of each videos (Sec. 3.2). Starting from the initial estimations, we further refine them by jointly optimizing calibration parameters meanwhile reconstructing dynamic NeRF (Sec. 3.3).  Figure 3 describes the overall pipeline of our approach.
@@ 3.1 Problem Setup and Calibration Preparation
We reconstruct dynamic NeRF from $N$ unsynchronized multi-view videos $\{V^{i}\}_{i=1}^{N}$ whose poses are not known. Each video $V^{i}$ contains $M^{i}$ image frames, $V^{i}=(I_{t}^{i};t\in[0,M^{i}-1])$, where $I_{t}$ is image of video frame at time $t$. We assume the dynamic scenes contain moving humans, but we do not assume that humans are the only dynamic objects. We estimate the time offset $\Delta t^{i}\in\mathbb{R}$ such that the timestamp $t$ for the $i$th video $V^{i}$ can be mapped to a synchronized global timestamp by adding time offset $t+\Delta t^{i}$. Furthermore, we aim to find camera poses in world coordinate, namely rotation $R^{i}_{t}$ and translation $\tau^{i}_{t}$ of each videos. [^1] We assume that camera intrinsics are known or already estimated with algorithms [12, 39]. Hereafter, we will use the term “calibration parameters” to refer to both time offsets and camera poses.
@@
$\mathbf{H}^{i}=\left(\bigoplus_{k=1}^{K}h_{k,t}^{i};t\in[0,M^{i}-1]\right),$ (1)
$h_{k,t}^{i}=\{\Phi_{k,t}^{i},\Theta_{k,t}^{i},\beta_{k}^{i},\Gamma_{k,t}^{i}\},$ (2)
where $K$ is the number of humans in the scene, $\oplus$ is concatenation operator, $h_{k,t}^{i}$ is state of the $k$th human at time $t$ which is composed of root orientation $\Phi_{k,t}^{i}\in\mathbb{R}^{3}$, body pose $\Theta_{k,t}^{i}\in\mathbb{R}^{22\times 3}$ that is modeled by relative 3D rotation of joints in axis-angle representation, position of root $\Gamma_{k,t}^{i}\in\mathbb{R}^{3}$, and human shape parameters $\beta_{k}^{i}\in\mathbb{R}^{16}$ which remains constant across whole time. Human motions $\mathbf{H}^{i}$s are extracted from each monocular video independently since videos are unsynchronized and we cannot leverage correspondence across different videos.
We can also obtain camera trajectory relative to human motion. First, we obtain camera trajectory with SLAM method [38] whose scale is ambiguous. Then, we recover scale with data-driven human motion priors during the optimization process following SLAHMR. For more details, we refer the readers to the original paper [44]. We demonstrate estimated human motion in Fig. 4. It is noteworthy that even the estimated human poses are not very accurate for some frames, we can robustly estimate calibration parameters with our initialization stage in Sec. 3.2.
Figure 4: Examples of estimated human motion on our dataset. While SLAHMR [44] can reconstruct valid human motion for most frames (left), it sometimes fails when there is an ambiguity due to the viewing angle (middle) or fast motion (right).
@@ 3.2 Initialization Stage
where $\mathbf{J}_{\text{canon}}\in\mathbb{R}^{22\times 3}$ is a 3D positions of human body joints at canonical space. Joint positions can be obtained with SMPL-H [34] model $\mathcal{S}$, which takes root rotation, body pose, and shape parameter as input. We set root rotation as $\mathbf{0}$ to get joint positions at canonical frame,

### v2

# Humans as a Calibration: Dynamic 3D Scene Reconstruction from Unsynchronized and Uncalibrated Videos

@@ Abstract
Recent works on dynamic 3D neural field reconstruction assume the input from synchronized multi-view videos whose poses are known. The input constraints are often not satisfied in real-world setups, making the approach impractical. We show that unsynchronized videos from unknown poses can generate dynamic neural fields as long as the videos capture human motion. Humans are one of the most common dynamic subjects captured in videos, and their shapes and poses can be estimated using state-of-the-art libraries. While noisy, the estimated human shape and pose parameters provide a decent initialization point to start the highly non-convex and under-constrained problem of training a consistent dynamic neural representation. Given the shape and pose parameters of humans in individual frames, we formulate methods to calculate the time offsets between videos, followed by camera pose estimations that analyze the 3D joint positions. Then, we train the dynamic neural fields employing multiresolution grids while we concurrently refine both time offsets and camera poses. The setup still involves optimizing many parameters; therefore, we introduce a robust progressive learning strategy to stabilize the process. Experiments show that our approach achieves accurate spatio-temporal calibration and high-quality scene reconstruction in challenging conditions. Project page
@@ 1 Introduction
Recent advances in neural radiance fields (NeRF) have been extended to dynamic scenes. However, even with a series of regularization techniques, obtaining a dynamic 3D volume with video inputs is a highly under-constrained problem and often suffers from instability and convergence problems. The claimed success of dynamic NeRF assumes unrealistic movements of a monocular camera that can emulate a multi-view setup [9] or perfectly calibrated and synchronized multi-view inputs. Such input constraints are hard to achieve in the real world. Time synchronization often relies on hard-wired devices [13, 10] or needs additional cues [6] (e.g., sound). Pose estimation methods struggle in scenes with textureless areas or repetitive structures, as shown in Fig. 2. Furthermore, it becomes more challenging if scenes are dynamic and videos are not synchronized.
@@
We aim to allow training dynamic NeRF from a set of casually captured real-world videos of a shared event, such as a sports game or a concert, which can only be collected later. In other words, we reconstruct photorealistic 4D scenes from unsynchronized multi-view videos with unknown camera poses. We need strong priors to tackle this challenging problem with high degrees of freedom. We focus on humans, one of the most common dynamic objects in scenes. Human pose estimation in computer vision has progressed rapidly and now achieves reliable performance even in general images or videos. We consider an estimated human parameter a robust mid-level representation to deduce the relationship between unsynchronized videos. Namely, we exploit human motion as a calibration pattern to estimate both time offsets and camera poses.
@@
We first find the initial time offsets and camera poses using the sequence of human shape and pose parameters estimated from individual videos. We can obtain a robust estimation of time offsets by enforcing global consistency in the sequence of motions with pairwise scores. Specifically, we consider the estimated parametric model of humans as a time series whose feature vector is the pose and shape parameters. We apply dynamic time warping (DTW) between every pair of videos to compute a constant offset to minimize the distance between 3D joint positions and record the associated cost value. The pairwise values are stored within a matrix, from which we find a global sequence alignment with the overall minimum costs. We estimate camera poses at coherent world coordinates from the alignment of human motions, by applying Procrustes analysis on aggregated 3D joint positions across overlapping frames.
@@
Even with slight misalignments of camera parameters, NeRF reconstruction degrades significantly. Therefore, we further refine the estimated time offsets and camera poses by jointly optimizing them with dynamic NeRF [8] training. We modify coarse-to-fine registration [21] to stabilize the training of the multiresolution grid-based representation. Furthermore, we propose a curriculum learning strategy that adds more variables as the optimization progresses. Optimization scheduling is critical for convergence in gradient-based optimization.
@@
We evaluate our approach in various challenging real-world datasets including CMU Panoptic Studio [13], Mobile-Stage [46], and EgoBody [51]. Our initialization stage with human motion alignment robustly estimates both time offsets and camera poses. Also, refinement with joint optimization achieves both highly accurate calibration and high-quality dynamic 3D scene reconstruction.
@@ 2 Related Works
Recent human pose and shape estimation methods leverage powerful data-driven prior from abundant datasets containing humans [12, 26, 1, 22, 10]. Most approaches recover human pose and shape by estimating parametric models such as SMPL [24]. From optimization-based methods [3, 35] to regression-based methods [14, 15, 17], early works faithfully recover 3D human pose and shape in camera coordinate. Recent advances [18, 40] take a step forward to estimate the global trajectories of humans, which is crucial for a complete understanding of human motion. Notably, SLAHMR [48] utilizes human motion priors to recover both real-world scaled global human motion and camera trajectory, whose scale is ambiguous with the SLAM pipeline alone. We note that monocular video human shape and motion estimation works have matured and generalize well to novel scenes. In this paper, we explore the potential of human motion as a robust mid-level representation for multiple camera calibration in both temporal and spatial domain.
@@
Camera calibration with human
While the majority of camera calibration methods exploit image features, there are several recent works that utilize humans as a calibration cue. For example, Patel and Black [34] propose a single image camera intrinsics estimation network trained on a dataset of images containing people. Ma et al. [25] propose a bundle adjustment method with Virtual Correspondences between image pair. Xu et al. [45] utilize center of bounding boxes of detected humans as keypoints for wide-baseline camera calibration. Lee et al. [19] estimate camera poses with 3D joints of moving humans. Recently, HSfM [30] leverages humans to recover the metric scale of camera and scene coordinates obtained from DUSt3R [42]. Our work shares the same key insights with this line of research; namely, humans provide powerful cues in challenging scenarios where traditional camera calibration method fails, such as scenes with repetitive structures, textureless scenes, or wide baseline setup. Also, utilizing humans allows us to recover the metric scale of reconstructed scenes and camera poses, which cannot be achieved with traditional SfM methods. However, previous methods only work for synchronized videos from static cameras [19] or static scenes [30, 45, 25]. Moreover, the estimated poses with humans are not accurate enough to directly reconstruct NeRF.
@@
The emergence of NeRF [29] and its extension to the temporal domain enables immersive experience in a dynamic world, which was previously limited to a small navigatable area and requires a complicated capturing system [4]. One of the most common approaches to reconstruct 4D scenes is reconstructing 3D scenes in a canonical frame and optimize a deformation field to warp them [31, 32, 36, 20, 23, 47]. However, as pointed out by Park et al. [33], it is hard to let the deformation network learn general dynamics in practice. On the other hand, recent grid-based methods [8, 5] simply represent 4D scenes by introducing an additional temporal domain without deformation field, which enables to represent general dynamic scenes easily. We exploit K-Planes [8] for our 4D scene representation due to its versatility.
However, all of the aforementioned dynamic NeRFs need synchronized videos and ground-truth camera poses. Recently, Sync-NeRF [16] aims to reconstruct dynamic NeRF from unsynchronized videos. They additionally optimize time offsets during the dynamic NeRF reconstruction. However, Sync-NeRF requires initial time offsets close to ground truth since it can only deal with small temporal perturbations and they also require known camera poses.
@@
Since iNeRF [49] has shown that one can optimize camera poses given pretrained NeRF with photometric loss, recent works attempt to train NeRF from unknown camera poses. They ease the burden of the strict requirements of NeRF (i.e., perfect camera poses) by jointly optimizing camera poses during NeRF training. Instead of naïve optimization [44], recent works propose coarse-to-fine optimization [21] and curriculum learning strategy that progressively adds camera parameters [11] for robust optimization. However, previous works only optimize camera poses with static NeRF. To the best of our knowledge, we are the first to reconstruct dynamic NeRF from unknown time offsets and camera poses. We also observe that progressive training is critical for robust optimization in dynamic NeRF scenarios. We propose a coarse-to-fine optimization technique for multiresolution grid representation and curriculum learning strategy. More importantly, all previous works need good initial points and only adjust small misalignments during joint optimization or only work for image sets with extremely small baselines (e.g., forward-facing capture [28]). We propose a novel way to obtain good initial points with humans, both for time offsets and camera poses.
@@ 3.1 Problem Setup and Calibration Preparation
We reconstruct dynamic NeRF from unsynchronized multi-view videos whose poses are not known. We assume that 1) dynamic scenes contain moving humans (we do not assume that humans are the only dynamic objects), 2) known intrinsics, and 3) known person correspondence across views when there are multiple humans. Note that we can handle an arbitrary number of humans in the scene. Also, we impose a minimal assumption on the camera parameters – we can handle moving cameras with different intrinsic or frame rates.
Each of $N$ multi-view input video $V^{i},i\in[1,N]$ contains $M^{i}$ image frames, $V^{i}=(I_{t}^{i};t\in[0,M^{i}))$, where $I_{t}$ is image frame at time $t$. We estimate time offset $\Delta t^{i}\in\mathbb{R}$ such that the timestamp $t$ for the $i$th video $V^{i}$ can be mapped to a synchronized global timestamp by adding time offset $t+\Delta t^{i}$. Furthermore, we aim to find camera poses in world coordinate, namely rotation $R^{i}_{t}$ and translation $\tau^{i}_{t}$ of each video. Hereafter, we will use the term “calibration parameters” to refer to both time offsets and camera poses.
@@
$\mathbf{H}^{i}{=}\left(\bigoplus_{k=1}^{K}h_{k,t}^{i};t\in[0,M^{i})\right),$ (1)
where $K$ is the number of humans in the scene, $\oplus$ is concatenation operator, $h_{k,t}^{i}{=}\{\Phi_{k,t}^{i},\Theta_{k,t}^{i},\beta_{k}^{i},\Gamma_{k,t}^{i}\}$ is state of the $k$th human at time $t$ which is composed of root orientation $\Phi_{k,t}^{i}\in\mathbb{R}^{3}$, body pose $\Theta_{k,t}^{i}\in\mathbb{R}^{22\times 3}$ that is modeled by relative 3D rotation of joints in axis-angle representation, position of root $\Gamma_{k,t}^{i}\in\mathbb{R}^{3}$, and human shape parameters $\beta_{k}^{i}\in\mathbb{R}^{16}$ which remains constant across whole time. Human motions $\mathbf{H}^{i}$s are extracted from each monocular video independently since videos are unsynchronized and we cannot leverage correspondence across different videos.
We can also obtain camera trajectory relative to human motion. First, we obtain camera trajectory with SLAM method [41] whose scale is ambiguous. Then, we recover scale with data-driven human motion priors during the optimization process following SLAHMR. For more details, we refer the readers to the original paper [48]. We demonstrate estimated human motion in Fig. 4. It is noteworthy that even if the estimated human poses are not very accurate for some frames, we can robustly estimate calibration parameters with our initialization stage in Sec. 3.2.
Figure 4: Examples of estimated human motion on our dataset. While SLAHMR [48] can reconstruct valid human motion for most frames (left), it sometimes fails when there is ambiguity due to the viewing angle (middle) or fast motion (right).
@@ 3.2 Initialization Stage
where $\mathbf{J}_{\text{canon}}\in\mathbb{R}^{22\times 3}$ is a 3D positions of human body joints relative to the root. Joint positions can be obtained with SMPL-H [37] model $\mathcal{S}$, which takes root rotation, body pose, and shape parameter as input. We set root rotation as $\mathbf{0}$ to get joint positions at canonical frame,